### v1

# Global bifurcation of localised 2D patterns emerging from spatial heterogeneity

Abstract
We present a general approach to prove the existence, both locally and globally in amplitude, of fully localised multi-dimensional patterns in partial differential equations containing a compact spatial heterogeneity. While one-dimensional localised patterns induced by spatial heterogeneities have been well-studied, proving the existence of fully localised patterns emerging from a Turing instability in higher dimensions remains a key open problem in pattern formation. In order to demonstrate the approach, we consider the two-dimensional Swift–Hohenberg equation, whose linear bifurcation parameter is perturbed by a radially-symmetric potential function. In this case, the trivial state is unstable in a compact neighbourhood of the origin and linearly stable outside. We prove the existence of local bifurcation branches of fully localised patterns, characterise their stability and bifurcation structure, and then rigorously continue solutions to large amplitude via analytic global bifurcation theory. Notably, the primary bifurcating branch in the Swift–Hohenberg equation alternates between an axisymmetric spot and a non-axisymmetric ‘dipole’ pattern, depending on the width of the spatial heterogeneity.
1 Introduction
In this paper we prove the global existence of fully localised two-dimensional patterns close to a pattern-forming or Turing instability. The key breakthrough is to use a compact region of spatial heterogeneity to induce a pattern-forming bifurcation away from the Turing point, where the linearised operator remains Fredholm. Although there are many proofs—particularly in the area of water waves—of global bifurcations of waves and patterns, the extension to fully localised multi-dimensional patterns remains elusive. As such, this work provides a significant novel direction for future research.
The emergence of patterns with spatially localised extent are important for understanding phenomena in a wide range of applications from the formation of tornadoes [53, 15], patches and fairy circles in dryland vegetation [12, 39, 33, 52], and spikes on the surface of a magnetic fluid [50, 59]; see [7] for a review of localised pattern formation. The key bifurcation of these localised patterns from quiescence is due to a pattern-forming instability [7]. While there is extensive theory for the emergence of one-dimensional localised patterns, to date there remains no existence proof for fully localised 2D patterns emerging from a Turing instability beyond the axisymmetric case. Fundamentally, this is due to the fact that the kernel of the linear operator at a 1D pattern-forming instability is finite dimensional but in 2D it is infinite-dimensional. We note that there have been recent advances in approximating these patterns [34, 35, 36], as well as in proving their existence at certain parameter regimes [61, 14, 51] via rigorous numerics.
The present work is inspired by the use of spatial heterogeneity to induce the emergence of localised patterns in experiments. Two examples are the localised temperature hotspots that can lead to the nucleation of tornadoes in numerical simulations [53, 15], and the local amplification of a magnetic field inducing axisymmetric spikes to form on the surface of a magnetic fluid [59]. The latter example directly informs our approach in this work. The experiments of Richter and Barashenkov [59] utilise a probe coil to destabilise the flat state in a compact region to generate an axisymmetric soliton (see [37] for a previous study of axisymmetric ferrofluid spikes in the absence of spatial heterogeneity). The induced spike persists following the removal of the probe coil, allowing the authors to study the behaviour of localised solutions to the spatially homogeneous problem. The Swift–Hohenberg equation we study in this work is taken to be a toy model for the probe coil problem, where we introduce a compactly-supported potential function to mimic the linear stability problem in the ferrofluid experiment. We note that there has been significant recent interest in the effects of spatial heterogeneity on pattern formation (see [42, 2, 26, 1, 27, 16, 28, 41, 40, 43] for stripe/target deformation, [5, 22, 23, 44, 45, 56] for pattern formation in parabolic equations, and [9, 48, 21] in hyperbolic equations); the present work fits well within this expanding literature, while also developing a new direction for the rigorous study of localised multi-dimensional patterns.
In this paper, we develop a new theory for the existence of large amplitude localised patterns with dihedral symmetry, where the spatial heterogeneity in the linear operator is linearly destabilising in a compact disc centred at the origin and stabilising everywhere else; see Figure 1. The key advantage of this approach is that for this type of spatial heterogeneity the essential spectrum is bounded away from the imaginary axis and allows the use of Fredholm operator theory, overcoming the fundamental difficulty in the spatially homogeneous case. One is then left with the problem of finding bifurcating point spectra; we show how this can be done where the linear operator is isotropic, though we believe that this requirement can be relaxed. Provided one can then locate bifurcating point spectra, one can then employ Crandall–Rabinowitz theory [19] and global bifurcation theory [57, 11, 18] to prove the existence of branches of large-amplitude, nonlinear fully localised patterns.
Figure 1: We consider a PDE system appended with a localised potential $V(\mathbf{x})$, causing the otherwise-stable trivial state to destabilise in a compact region and inducing the emergence of localised patterns.
The Crandall–Rabinowitz theorem allows one to prove the existence of small amplitude solutions bifurcating from a given base state while analytic global bifurcation theory—originally due to Dancer [20] and later developed by Buffoni and Toland [11] and Constantin, Strauss and Vărvărucă [18]—allows one to continue small amplitude solutions to large amplitude. The Crandall–Rabinowitz theorem is typically applied to bounded domains or spatially periodic solutions; while localised solutions have been considered in some cases [64, 60], the application of Crandall–Rabinowitz theory is more difficult due to a lack of standard embedding and compactness properties in infinitely-extended spatial domains. While analytic global bifurcation theory has become invaluable in proving the existence of large amplitude water waves, including some recent applications to solitary waves (see [31, Section 5.6] and references therein), very few results exist in the pattern formation literature (see for instance [54, 24, 4, 25] which use the global bifurcation theory of Rabinowitz [57], and the more recent work of [10, 13] which use analytic global bifurcation theory). A key difficulty with applying the theory of Buffoni and Toland [11] to localised solutions is a possible lack of compactness of solution sets. While in bounded domains (or, equivalently, for periodic solutions) local compactness of the solution set is guaranteed by a combination of Schauder estimates and Rellich–Kondrachev-type embeddings, these results no longer hold in infinitely extended domains. A crucial change in this approach was provided by Chen, Walsh and Wheeler in [17], where the authors adapted the analytic global continuation theorem of Buffoni and Toland so that local compactness was no longer an assumption, but rather a property of the solution set that could fail. To the best of the authors’ knowledge, there are no global bifurcation results for localised patterns in pattern-forming systems, and so this paper presents a new application of the theory for 2D fully localised patterns with dihedral symmetry.
Our approach is to consider the following semi-linear PDE system. Fix $d\in\mathbb{N}$ with $d>1$, $N\in\mathbb{N}$, and consider
$\partial_{t}u=\mathcal{L}_{\varepsilon}(\Delta)u+V_{\varepsilon}(|\mathbf{x}|)\,u+\mathcal{N}_{\varepsilon}(u),\qquad\qquad u\in H^{2m}(\mathbb{R}^{d};\mathbb{R}^{N}),\qquad\varepsilon\in\Lambda\subseteq\mathbb{R},$ (1.1)
with $\mathcal{L}_{\varepsilon}(\Delta):H^{2m}(\mathbb{R}^{d};\mathbb{R}^{N})\to L^{2}(\mathbb{R}^{d};\mathbb{R}^{N})$ an isotropic linear operator (where $\Delta$ denotes the $d$-dimensional Laplace operator and $m>d/4$), $V_{\varepsilon}$ a localised radially-symmetric potential, and $\mathcal{N}_{\varepsilon}$ a real-analytic function with $\mathcal{N}_{\varepsilon}(0)=\mathcal{N}_{\varepsilon}^{\prime}(0)=0$ for all $\varepsilon\in\Lambda$. We additionally assume that $\mathcal{L}_{\varepsilon},V_{\varepsilon},\mathcal{N}_{\varepsilon}$ are each analytic in $\varepsilon\in\Lambda$, and that the following hypotheses are satisfied.
Hypotheses.
  1. (i)
    $\mathcal{L}_{\varepsilon}(-|\mathbf{k}|^{2})<0$ for all $\mathbf{k}\in\mathbb{R}^{d}$, $\varepsilon\in\Lambda$ ;
  2. (ii)
    there exists some $\varepsilon_{0}\in\Lambda$ such that $\mathrm{ker}\,(\mathcal{L}_{\varepsilon_{0}}(\Delta)+V_{\varepsilon_{0}})=\mathrm{span}\,\{v_{0}\}$ for some $v_{0}\in X\subseteq H^{2m}(\mathbb{R}^{d};\mathbb{R}^{N})$ [^1] One must restrict to some subspace $X$ of $H^{2m}(\mathbb{R}^{d};\mathbb{R}^{N})$ in order to remove innate symmetries of the problem and obtain a one-dimensional kernel.; and
  3. (iii)
    $\partial_{\varepsilon}(\mathcal{L}_{\varepsilon}(\Delta)+V_{\varepsilon})_{\varepsilon=\varepsilon_{0}}v_{0}\notin\mathrm{Im}(\mathcal{L}_{\varepsilon_{0}}(\Delta)+V_{\varepsilon_{0}})$.
Hypothesis (i) supposes that the operator $\mathcal{L}_{\varepsilon}(\Delta)$ and the linearisation of (1.1) about some $u=\phi\in H^{2m}(\mathbb{R}^{d};\mathbb{R}^{N})$ are Fredholm operators with index zero. In particular, the essential spectrum of each operator is bounded away from zero and so any changes in the linear stability of solutions to (1.1) is driven entirely by eigenvalues crossing the imaginary axis. Hypothesis (ii) supposes that, for some parameter value $\varepsilon=\varepsilon_{0}$, there is one eigenvalue lying on the imaginary axis with a single corresponding eigenfunction $v_{0}$. Finally, hypothesis (iii) supposes that the equation (1.1) has nondegenerate dependence on the parameter $\varepsilon$ at $\varepsilon=\varepsilon_{0}$.
These hypotheses together tell us that the change in stability at $\varepsilon=\varepsilon_{0}$ corresponds to a transverse intersection between the trivial state $u\equiv 0$ and some non-trivial bifurcating solution $u=u_{*}(\mathbf{x})$. We are then able to apply local and global bifurcation results to prove the existence of non-trivial fully localised patterns along a continuous bifurcation curve
$\mathcal{C}=\{(\varepsilon,u)=(\mathcal{E}(s),v(s)):\;s\in\mathbb{R}\}.$
Here, $\mathcal{E}(0)=\varepsilon_{0}$, $v(0)=0$, $v^{\prime}(0)=v_{0}$, and the curve $\mathcal{C}$ possesses a local analytic parametrisation around each of its points. The local and global bifurcation results are proven in an open subspace $(\varepsilon,u)\in U\subseteq\Lambda\times H^{2m}(\mathbb{R}^{d};\mathbb{R}^{N})$, and so a solution $u=v(s)$ of (1.1) will remain spatially localised along the curve $\mathcal{C}$.
To illustrate our method, we consider the two-dimensional Swift–Hohenberg equation given by
$\partial_{t}u=-[(1+\Delta)^{2}+\varepsilon^{2}]u+2\,\varepsilon^{2}\,V(|\mathbf{x}|)\,u+f(u),\qquad\quad\text{with}\qquad V(r)=\begin{cases}1,&r<R,\\ 0,&r>R.\end{cases}$ (1.2)
where $u=u(t,\mathbf{x})\in\mathbb{R}$, $\mathbf{x}\in\mathbb{R}^{2}$, $R>0$ denotes the width of the potential $V$, and $f$ is real-analytic. This equation naturally destabilises the trivial state in a finite circle of radius $R$ while outside this region the trivial state is linearly stable. We seek steady solutions $u=u(\mathbf{x})$ in $H^{4}_{\mathrm{e}}$—the subspace of functions in $H^{4}(\mathbb{R}^{2})$ which are also even in $y$, i.e. u(x,y)=u(x,y)u(x,y)=u(x,-y)—where we solve the steady equation
$0=-[(1+\Delta)^{2}+\varepsilon^{2}]u+2\,\varepsilon^{2}\,V(|\mathbf{x}|)\,u+f(u),$ (1.3)
posed in the analogous even subspace $L^{2}_{\mathrm{e}}$ of $L^{2}(\mathbb{R}^{2})$. We categorise possible solutions to (1.3) by introducing the following rotational subspaces
[…]
for any $k\in\mathbb{N}$ and $m\in\mathbb{N}_{0}$ (see Remark 1.1 for the case when $k=0$). We note that we can equivalently define the spaces $X^{m}_{k}$ as the intersections of $H^{m}(\mathbb{R}^{2};\mathbb{R})$ Sobolev spaces and $\mathbb{D}_{k}$ dihedral symmetry spaces, where we fix the dihedral line of reflection symmetry to be the x-axis.
Remark 1.1.
We define, with a slight abuse of notation, the space $X^{m}_{0}$ to be the subspace of $H^{m}(\mathbb{R}^{2};\mathbb{R})$ containing axisymmetric functions $u(\mathbf{x})=u(|\mathbf{x}|)$; we likewise define $\mathbb{D}_{0}$ to be the set of axisymmetric functions in this notation. We also note that $X^{m}_{1}$ is the entire space $H^{m}_{\mathrm{e}}$ but we maintain the vacuous notation $X^{m}_{1}$ and $\mathbb{D}_{1}$ in order to be consistent with other values of $k\in\mathbb{N}_{0}$.
Remark 1.2.
The rotational subspaces form chains of subspaces
$X^{m+n}_{k}\subseteq X^{m}_{k},\qquad\qquad\qquad X^{m}_{n\,k}\subseteq X^{m}_{k}$
for any $k,m,n\in\mathbb{N}_{0}$. Furthermore, the dihedral space $\mathbb{D}_{k}$ represents an invariant subspace of the equation (1.2) for each $k\in\mathbb{N}_{0}$.
For our particular choice of spatial heterogeneity, we are able to derive explicit characterisations for the point spectrum of the linearisation of (1.2) about the trivial solution. In particular, we determine that eigenvalues $\lambda=\varepsilon^{2}\widetilde{\lambda}$ solve the implicit relation
$F_{k}(R,\varepsilon,\tilde{\lambda})=\mathrm{Re}\left[\mathrm{e}^{\mathrm{i}\sin^{-1}(\widetilde{\lambda})}\,W[J_{k}(\alpha_{+}r),H^{(1)}_{k}(\beta r)](R)\,W[J_{k}(\alpha_{-}r),\overline{H^{(1)}_{k}(\beta r)}](R)\right]=0$
for some $k\in\mathbb{N}_{0}$ and fixed $\varepsilon,R>0$. Here, $\alpha_{\pm}:=\sqrt{1\pm\varepsilon\,\sqrt{1-\widetilde{\lambda}}}$, $\beta:=\sqrt{1+\mathrm{i}\,\varepsilon\,\sqrt{1+\widetilde{\lambda}}}$; $J_{k},H_{k}^{(1)}$ are the $k$’th order Bessel functions of the first and third kind; and $W[\cdot,\cdot]$ denotes the weighted Wronskian function given by
$W[u,v](r):=r\left(u(r)\,v^{\prime}(r)-u^{\prime}(r)\,v(r)\right).$
We numerically observe that there are infinitely many bifurcation points $\varepsilon=\varepsilon_{k,n}$ for $k\in\mathbb{N}_{0}$ and $n\in\mathbb{N}$, and that the parameter $R$ (corresponding to the width of the potential $V$) can be chosen such that the above hypotheses are satisfied.
Figure 2: (a) A schematic bifurcation diagram showing different types of localised dihedral patterns bifurcating off the trivial state. (b) A local bifurcation curve ($b_{1}$) can be extended to a global bifurcation curve ($b_{2}$) using analytic bifurcation theory. In particular, the analytic curves that emanate from the local bifurcation either: $(i)$ blow up in their norm, $(ii)$ collide with the boundary $\varepsilon=0$, $(iii)$ form closed loops, or $(iv)$ lose compactness.
Our aim is then to prove the schematic picture shown in Figure 2(a). We note that we prove the existence of fully localised patterns with $\mathbb{D}_{k}$ symmetry for any $k\in\mathbb{N}_{0}$, and not just the $\mathbb{D}_{0}$, $\mathbb{D}_{1}$ and $\mathbb{D}_{6}$ localised patterns shown in Figure 2(a). We employ analytic bifurcation theory in order to prove the existence of the bifurcating localised patterns, as illustrated in Figure 2(b); we first prove the existence of a unique non-trivial solution branch in a neighbourhood of the bifurcation point (Figure 2($b_{1}$)), which we call the local branch, before extending solutions to large amplitude (Figure 2($b_{2}$)), which we call the global branch. The global solution branch is either unbounded, approaches the $\varepsilon=0$ axis, forms a closed loop, or loses compactness, as shown in Figure 2(ii)-(iviv).
We present two main theorems for the existence and stability of fully localised 2D pattern solutions to (1.2). We begin by informally stating our existence result, where we refer to Theorems 3.3 and 3.9 for the precise statements.
Result 1 (Existence).
Fix $k\in\mathbb{N}_{0}$ and $R>0$. For any $\varepsilon_{0}$ that is a simple root of $F_{k}(R,\varepsilon_{0},0)=0$ (and is not a root for any other $k\in\mathbb{N}_{0}$), there exists a localised solution $u\in X^{4}_{k}$ which bifurcates from $(\varepsilon,u)=(\varepsilon_{0},0)$ and continues to a global solution branch.
Amongst the many bifurcating solutions we are considering, we are particularly interested in the initial destabilisation of the trivial state. We call this point the primary bifurcation, corresponding to the the minimum value of $\varepsilon$ that solves $F_{k}(R,\varepsilon,0)=0$ for some $k\in\mathbb{N}_{0}$. We will presently observe that the primary branch consists of qualitatively different solutions for different parameter values, alternating between axisymmetric and non-axisymmetric profiles as $R$ increases.
We prove the linear (in)stability of localised solutions to (1.2) near their bifurcation, where we note qualitative differences between the primary and non-primary bifurcation points. We again present an informal version of our stability result here and refer the reader to Lemma 3.6 for more details.
Result 2 (Stability).
Near the primary bifurcation point, the bifurcating solution is linearly stable if $\varepsilon>\varepsilon_{0}$ and unstable if $\varepsilon<\varepsilon_{0}$. Subsequent bifurcation points only yield locally unstable branches.
The paper is outlined as follows. In §2 we characterise the linear spectrum of (1.2) and prove certain properties required for our bifurcation results. We then present our local and global bifurcation theorems in §3, where we also prove the type of bifurcation and local stability of the bifurcating solutions. We finally conclude and discuss open problems in §4.
2 Spectral properties of the linear operator
We begin by considering the linear stability of solutions to (1.2). In particular, we analyse the stability of a general solution $\phi\in H^{4}_{\mathrm{e}}$—including the trivial solution $\phi\equiv 0$—with respect to continuous perturbations, thus characterising the essential spectrum of the linear problem (§2.1). We then characterise the point spectrum of the linear problem by constructing the eigenvalues and critical eigenfunctions that determine the stability of the trivial solution (§2.2).
We consider the linear operator $\mathcal{L}$, given by
$\mathcal{L}u=-(1+\Delta)^{2}u-\varepsilon^{2}\,u+2\,\varepsilon^{2}\,V(|\mathbf{x}|)\,u$ (2.1)
as well as the linearisation $\mathcal{L}_{\phi}$ of (1.2) about a function $\phi\in H^{4}_{\mathrm{e}}$, given by
$\mathcal{L}_{\phi}u=-(1+\Delta)^{2}u-\varepsilon^{2}\,u+2\,\varepsilon^{2}\,V(|\mathbf{x}|)\,u+f^{\prime}(\phi(\mathbf{x}))\,u.$ (2.2)
2.1 Essential spectrum
We note that (2.1) is a special case of (2.2) with $\phi\equiv 0$, and so we present the following result for the general linear operator $\mathcal{L}_{\phi}$ with $\phi\in H^{4}_{\mathrm{e}}$.
Lemma 2.1.
Fix $\varepsilon,R>0$ and suppose that $\phi\in H^{4}_{\mathrm{e}}$. The linear operator $\mathcal{L}_{\phi}:H^{4}_{\mathrm{e}}\to L^{2}_{\mathrm{e}}$ given by (2.2) is a Fredholm operator with index 0. In particular, $\sigma_{\mathrm{ess}}(\mathcal{L}_{\phi})=(-\infty,-\varepsilon^{2}]$, where $\sigma_{\mathrm{ess}}(\mathcal{L}_{\phi})$ denotes the essential spectrum of $\mathcal{L}_{\phi}$.
Proof.
We first note that the operator $\mathcal{L}_{\phi}$ can be decomposed into three operators $\mathcal{T},\mathcal{K}_{V},\mathcal{K}_{\phi}:H^{4}_{\mathrm{e}}\to L^{2}_{\mathrm{e}}$ by $\mathcal{L}_{\phi}=\mathcal{T}+\mathcal{K}_{V}+\mathcal{K}_{\phi}$, where
$\mathcal{T}u=-[(1+\Delta)^{2}+\varepsilon^{2}]\,u,\qquad\qquad\mathcal{K}_{V}\,u=2\,\varepsilon^{2}\,V(|\mathbf{x}|)\,u,\qquad\qquad\mathcal{K}_{\phi}u=f^{\prime}(\phi(\mathbf{x}))\,u.$
Using the fact that $V(|\mathbf{x}|)\,u\in H^{4}(B_{C})$, where $B_{C}$ denotes the open ball of radius $C$ about the origin, and applying the Rellich–Kondrachov theorem, we conclude that $\mathcal{K}_{V}:H^{4}_{\mathrm{e}}\to L^{2}_{\mathrm{e}}$ is a compact operator. Likewise, since $f$ is real analytic and $\phi\in H^{4}_{\mathrm{e}}$, we conclude that $f^{\prime}(\phi)\in H^{4}_{\mathrm{e}}$ and thus, by Lemma B.1, $K_{\phi}:H^{4}_{\mathrm{e}}\to L^{2}_{\mathrm{e}}$ is also compact. It follows that $\mathcal{L}_{\phi}$ is Fredholm if and only if $\mathcal{T}$ is Fredholm. Using the two-dimensional Fourier transform, we note that the operator $\mathcal{T}$ is unitarily equivalent to multiplication by its symbol
$\widehat{\mathcal{T}}=-(1-|\mathbf{k}|^{2})^{2}-\varepsilon^{2},\qquad\qquad\mathbf{k}\in\mathbb{R}^{2}$
and so we conclude that the operators $\mathcal{T}$ and $\mathcal{L}_{\phi}$ are Fredholm with index 0 for all $\varepsilon>0$. Since the essential spectrum is preserved under compact perturbations, it follows that $\sigma_{\mathrm{ess}}(\mathcal{L}_{\phi})=\sigma_{\mathrm{ess}}(\widehat{\mathcal{T}})=(-\infty,-\varepsilon^{2}]$. ∎
Remark 2.2.
The linear operator $\mathcal{L}_{2}:H^{2}_{\mathrm{e}}\times H^{2}_{\mathrm{e}}\to L^{2}_{\mathrm{e}}\times L^{2}_{\mathrm{e}}$, given by
$\begin{split}\mathcal{L}_{2}\mathbf{u}={}&\begin{pmatrix}\Delta+1&-1\\ \varepsilon^{2}-2\,\varepsilon^{2}\,V(|\mathbf{x}|)&\Delta+1\end{pmatrix}\mathbf{u},\end{split}$
is isomorphic to $\mathcal{L}$ by the mapping $\mathcal{S}:H^{4}_{\mathrm{e}}\to H^{2}_{\mathrm{e}}\times H^{2}_{\mathrm{e}},\,u\mapsto(u,(1+\Delta)u)$, and thus is also a Fredholm operator with index zero. One would thus expect the results presented herein to extend to reaction-diffusion systems without significant changes.
2.2 Point spectrum
In the previous section, we determined that the linear operator $\mathcal{L}:H^{4}_{\mathrm{e}}\to L^{2}_{\mathrm{e}}$ is a Fredholm operator with index 0 for $\varepsilon>0$ and, in particular, that $\sigma_{\mathrm{ess}}(\mathcal{L})=(-\infty,-\varepsilon^{2}]$. This tells us that the essential spectrum of $\mathcal{L}$ is bounded away from 0, and so any changes in the linear stability of $\mathcal{L}$ must be driven by eigenvalues crossing the imaginary axis.
We now characterise the point spectrum of $\mathcal{L}$ by considering the linear eigenvalue problem $\lambda u=\mathcal{L}\,u$, where $u\in H^{4}_{\mathrm{e}}$ is an eigenfunction of $\mathcal{L}$ with associated eigenvalue $\lambda\in\mathbb{C}$. For notational simplicity, we introduce some $\tilde{\lambda}\in\mathbb{C}$, with $\lambda=\tilde{\lambda}\,\varepsilon^{2}$, and consider the scaled eigenvalue problem
$\varepsilon^{2}\,\tilde{\lambda}\,u=-(1+\Delta)^{2}u-\varepsilon^{2}u+2\,\varepsilon^{2}\,V(|\mathbf{x}|)\,u.$ (2.3)
Lemma 2.3.
Fix $R>0$ and $\varepsilon>0$. The eigenvalue problem (2.3) possesses a separable solution $u=u_{k}\in H^{4}_{\mathrm{e}}$ given by $u_{k}(r\cos\theta,r\sin\theta)=v_{k}(r)\,\cos(k\theta)$ with
[…]
where $\alpha_{\pm}:=\sqrt{1\pm\varepsilon\,\sqrt{1-\tilde{\lambda}}}$, $\beta:=\sqrt{1+\mathrm{i}\,\varepsilon\,\sqrt{1+\tilde{\lambda}}}$; $J_{k},H_{k}^{(1)}$ are the $k$’th order Bessel functions of the first and third kind; $\psi$ is given by $\psi=\frac{1}{2}\sin^{-1}(\tilde{\lambda})+\frac{\pi}{4}$; $\phi$ is given by
$\mathrm{e}^{2\mathrm{i}\phi}=\mathrm{i}\,\mathrm{e}^{-\mathrm{i}\sin^{-1}(\tilde{\lambda})}\frac{W[J_{k}(\alpha_{+}r),\overline{H^{(1)}_{k}(\beta r)}](R)\,H^{(1)}_{k}(\beta R)}{W[J_{k}(\alpha_{+}r),H^{(1)}_{k}(\beta r)](R)\,\overline{H^{(1)}_{k}(\beta R)}};$
$W[\cdot,\cdot]$ denotes the weighted Wronskian function given by
$W[u,v](r):=r\left(u(r)\,v^{\prime}(r)-u^{\prime}(r)\,v(r)\right);$
and $\tilde{\lambda}=\tilde{\lambda}(R,\varepsilon)$ satisfies
$F_{k}(R,\varepsilon,\tilde{\lambda})=\mathrm{Re}\left[\mathrm{e}^{\mathrm{i}\sin^{-1}(\tilde{\lambda})}\,W[J_{k}(\alpha_{+}r),H^{(1)}_{k}(\beta r)](R)\,W[J_{k}(\alpha_{-}r),\overline{H^{(1)}_{k}(\beta r)}](R)\right]=0$
with $\tilde{\lambda}\notin(-\infty,-1]$.
Figure 3: Radial profiles (resp. planar profiles) of linear eigenfunctions $v_{k,n}(r)$ ($u_{k,n}(\mathbf{x})$) are plotted for $R=8$ and (a) $\varepsilon=\varepsilon_{0,1}$, (b) $\varepsilon=\varepsilon_{1,1}$, and (c) $\varepsilon=\varepsilon_{6,1}$.
The proof of this lemma is given in Appendix B; we illustrate examples of the linear eigenfunctions $u=u_{k,n}$ in Figure 3. We highlight the respective $\mathbb{D}_{0}$ and $\mathbb{D}_{1}$ eigenfunctions $u_{0,1}$ and $u_{1,1}$, plotted in Figures 3 (a) and (b), which we refer to as the ‘axisymmetric spot’ and ‘dipole’ patterns. The equation $F_{k}(R,\varepsilon,\tilde{\lambda})=0$ implicitly defines eigenvalues $\tilde{\lambda}=\tilde{\lambda}(R,\varepsilon)$ for fixed $R,\varepsilon>0$, and the associated bifurcation equation $F_{k}(R,\varepsilon,0)=0$ equivalently defines bifurcation points $\varepsilon=\varepsilon(R)$ for a fixed $R>0$. While we do not obtain an explicit form for the eigenvalues $\tilde{\lambda}$, we are able prove the following qualitative properties.
Lemma 2.4.
An eigenvalue $\tilde{\lambda}\in\mathbb{C}$ of (2.3) satisfies the following properties:
  1. (i)
    $\tilde{\lambda}\in\mathbb{R}$ and $1<\tilde{\lambda}\leq 1$ for all $R,\varepsilon>0$;
  2. (ii)
    $\tilde{\lambda}\to-1$ as $R\to 0$ for any fixed $\varepsilon>0$; and
  3. (iii)
    for each $k\in\mathbb{N}_{0}$ and sufficiently large $R$ with $\varepsilon=\tilde{\varepsilon}\,R^{-1}$, there are countably many eigenvalues $\{\tilde{\lambda}_{k,n}\}_{n=1}^{\infty}$ which solve the implicit equation
    $\tilde{\varepsilon}=\frac{2\sin^{-1}(\tilde{\lambda}_{k,n})+(2n-1)\pi}{2\,\sqrt{1-\tilde{\lambda}_{k,n}}}.$
    The eigenvalue $\tilde{\lambda}_{k,n}$ is a monotone increasing function of $\tilde{\varepsilon}$ for each $n\in\mathbb{N}$.
Proof.
We begin by deriving simple a priori estimates for an eigenvalue of (2.3). The operator $\mathcal{L}$ is symmetric with respect to the complex $L^{2}$-inner product and so it follows that $\tilde{\lambda}\in\mathbb{R}$. Assuming that $u$ satisfies (2.3) for some $\tilde{\lambda}\in\mathbb{R}$, we take the $L^{2}$-inner product of (2.3) with $u$ and, applying Green’s identities, obtain
$\tilde{\lambda}=2\,\frac{\|V\,u\|^{2}}{\|u\|^{2}}-\frac{\|(1+\Delta)u\|^{2}}{\varepsilon^{2}\,\|u\|^{2}}-1.$ (2.4)
It follows from the simple estimate $\|V\,u\|^{2}\leq\|u\|^{2}$ that $\tilde{\lambda}\leq 1$, where we recall that $\tilde{\lambda}>-1$ follows from Lemma 2.3. Furthermore, we note that
$\|V\,u\|^{2}=\int_{B_{R}}|u(\mathbf{x})|^{2}\,\mathrm{d}\mathbf{x}\to 0\qquad\text{as $R\to 0$}$
and so there exists some $R_{0}(\varepsilon)>0$ such that $\tilde{\lambda}<0$ for all $R\leq R_{0}$ and fixed $\varepsilon>0$.
We recall from Lemma 2.3 that an eigenvalue $\tilde{\lambda}$ of (2.3) must satisfy
$F_{k}(R,\varepsilon,\tilde{\lambda})=\mathrm{Re}\left[\mathrm{e}^{\mathrm{i}\sin^{-1}(\tilde{\lambda})}\,W[J_{k}(\alpha_{+}r),H^{(1)}_{k}(\beta r)](R)\,W[J_{k}(\alpha_{-}r),\overline{H^{(1)}_{k}(\beta r)}](R)\right]=0$
for some $k\in\mathbb{N}_{0}$, and we note the following asymptotic expansions:
  1. (a)
    for fixed $\varepsilon>0$
    $F_{k}(R,\varepsilon,\tilde{\lambda})=\frac{4}{\pi^{2}}\left(\frac{\alpha_{+}\,\alpha_{-}}{|\beta|^{2}}\right)^{k}\,\sqrt{1-\tilde{\lambda}^{2}}+\mathcal{O}(R)$
    as $R\to 0$; and
  2. (b)
    for fixed $\varepsilon=\tilde{\varepsilon}\,R^{-1}$ with fixed $\tilde{\varepsilon}>0$
    $F_{k}(R,\tilde{\varepsilon}\,R^{-1},\tilde{\lambda})=\frac{4}{\pi^{2}}\,\mathrm{e}^{-\sqrt{1+\tilde{\lambda}}\,\tilde{\varepsilon}}\,\cos\left(\sin^{-1}(\tilde{\lambda})-\sqrt{1-\tilde{\lambda}}\,\tilde{\varepsilon}\right)+\mathcal{O}(R^{-1})$
    as $R\to\infty$.
It follows that $\tilde{\lambda}^{2}\to 1$ as $R\to 0$ for fixed $\varepsilon>0$, where the a priori estimates tell us that in fact $\tilde{\lambda}\to-1$.
In the case when $\varepsilon=\tilde{\varepsilon}\,R^{-1}$ and $R\to\infty$, the eigenvalue $\tilde{\lambda}$ satisfies
$\tilde{\varepsilon}=g_{k,n}(\tilde{\lambda}):=\frac{2\,\sin^{-1}(\tilde{\lambda})+(2n-1)\pi}{2\,\sqrt{1-\tilde{\lambda}}}$
for any $n\in\mathbb{N}$. We note that $\sin^{-1}(\tilde{\lambda})>-\frac{\pi}{2}$ for $\tilde{\lambda}\in(-1,1)$, and so $g_{k,n}(\tilde{\lambda})>0$ for each $n\in\mathbb{N}$. Furthermore,
$g_{k,n}^{\prime}(\tilde{\lambda})=\frac{1}{\sqrt{1-\tilde{\lambda}^{2}}\,\sqrt{1-\tilde{\lambda}}}+\frac{1}{2(1-\tilde{\lambda})}\,g_{k,n}(\tilde{\lambda})>\frac{1}{\sqrt{1-\tilde{\lambda}^{2}}\,\sqrt{1-\tilde{\lambda}}}>0$
[…]
Figure 4: (a) The most unstable eigenvalues $\lambda$ of $\mathcal{L}$ for $k=0,1,2,3$ are numerically computed and plotted for $R=5$ and $\varepsilon>0$; (b) - (e) a cartoon of the spectrum of the linear operator $\mathcal{L}$ is presented for critical values of $\varepsilon$, where only the essential spectrum and the four most unstable eigenvalues are shown.
Property $(iii)$ in Lemma 2.4 states that, for $R$ sufficiently large, the eigenvalues of (3.3) are monotone increasing in $\varepsilon$; we also numerically observe this behaviour for moderate values of $R$, as shown in Figure 4(a). We note that, while $\tilde{\lambda}_{k,n}$ appears independent of the value of $k\in\mathbb{N}_{0}$ in the asymptotic regime, the eigenvalues will vary in both $n$ and $k$ for moderate values of $R$. The qualitative behaviour of the spectrum of $\mathcal{L}$ is illustrated in Figure 4(b-e), where the essential spectrum (given by $\sigma_{\mathrm{ess}}(\mathcal{L})=(-\infty,-\varepsilon^{2}]$) moves to the left as $\varepsilon$ increases and point spectra move to the right, causing additional bifurcations as each eigenvalue crosses the imaginary axis.
We introduce the label $\varepsilon_{k,n}$ to be the value of $\varepsilon$ where $\tilde{\lambda}_{k,n}=0$. For sufficiently large values of $R$, we obtain that $\varepsilon_{k,n}=\frac{(2n-1)\pi}{2\,R}$ for all $k\in\mathbb{N}_{0}$ and $n\in\mathbb{N}$; we again numerically observe that each $\varepsilon_{k,n}$ persists for moderate values of $R$ and varies in both $n$ and $k$, as shown in Figure 5. In particular, we plot in Figure 5(a) the implicit curves for $F_{k}(R,\varepsilon,0)=0$ and observe that, while each implicit curve converges to a monotonic curve $\varepsilon_{k,n}=\frac{(2n-1)\pi}{2\,R}$ as $R\to\infty$, the curves exhibit greater oscillations for moderate values of $R$.
In order to use Crandall–Rabinowitz bifurcation theory, we require that the kernel of $\mathcal{L}$ is one dimensional at the bifurcation point. This is equivalent to saying that, for a given value of $R$, the curves in Figure 5(a-b) do not intersect one another for all $k\in\mathbb{N}_{0}$. While it may be possible to prove this holds for certain parameter regimes, we instead present the following hypothesis which we motivate by Figure 5.
Hypothesis 1.
We suppose that, for our choice of $R>0$ and $k\in\mathbb{N}_{0}$, the kernel of $\mathcal{L}:H^{4}_{\mathrm{e}}\to L^{2}_{\mathrm{e}}$ at $\varepsilon=\varepsilon_{k,n}$ is entirely spanned by the eigenfunction $u_{k,n}\in X^{4}_{k}$ given in Lemma 2.3.
We present the following analytic predictions based on Lemma 2.4 and Figure 5, which we will verify numerically in Section 3.1. We first note that as $R$ passes from $R=2.6$ to $R=2.8$, the primary bifurcating eigenfunction alternates from the axisymmetric spot $u_{0,1}$ to the dipole pattern $u_{1,1}$, plotted in Figures 3(a) and (b), respectively. Furthermore, we note that Figure 5(b) suggests that one can choose a value of $R$ such that at a given bifurcation point $\varepsilon=\varepsilon_{k,n}$ the linear operator $\mathcal{L}$ has a one-dimensional kernel; in particular, we predict that $u_{6,1}$ (plotted in Figure 3(c)) appears to satisfy this condition for $R=7.5$, as shown in Figure 5(d).
Figure 5: Implicit curves of $F_{k}(R,\varepsilon,0)=0$ for (a) $k=0,1,\dots,5$ and (b) $k=0,1,\dots,150$. The label $\mathbb{D}_{k}$ indicates the smallest dihedral group in which the linear eigenfunction $u=u_{k,n}$ lies, and the dotted black lines indicate the asymptotic curve $\varepsilon=\frac{(2n-1)\pi}{2\,R}$. (c) As $R$ passes between $2.6$ and $2.8$, the primary bifurcation swaps from the $\mathbb{D}_{0}$ pattern to the $\mathbb{D}_{1}$ pattern. (d) At $R=7.5$ the $\mathbb{D}_{6}$ pattern is an isolated root of $F_{k}(R,\varepsilon,0)=0$.
3 Bifurcation results
We now prove the existence of fully localised patterns bifurcating from trivial state. We formulate (1.2) as the bifurcation problem
$0=\mathscr{F}(\varepsilon,u):=-(1+\Delta)^{2}u-\varepsilon^{2}u+2\,\varepsilon^{2}\,V(|\mathbf{x}|)\,u+f(u)$ (3.1)
and prove the existence of a non-trivial solution curve $(\varepsilon,u)=(\mathcal{E}(s),v(s))$ parametrised by some $s\in\mathbb{R}$, first in a local neighbourhood of the bifurcation point (with $|s|\ll 1$) in §3.1, before extending to a global solution curve of large amplitude patterns (with $s\in\mathbb{R}$) in §3.2.
3.1 Local bifurcation
We begin by proving the existence of a unique bifurcation branch in a neighbourhood of the point $(\varepsilon,u)=(\varepsilon_{k,n},0)$, which we call the local bifurcation branch. We will use an analytic version of the Crandall–Rabinowitz theorem (cf. Buffoni and Toland [11, Theorem 8.3.1]), which we state as Theorem A.1 in the appendix. In order to prove that a local bifurcation occurs, we need to verify the following conditions:
  1. $(i)$
    $\mathcal{L}:H^{4}_{\mathrm{e}}\to L^{2}_{\mathrm{e}}$ is a Fredholm operator of index zero at $\varepsilon=\varepsilon_{k,n}$,
  2. $(ii)$
    $\mathrm{ker}(\mathcal{L})=\mathrm{span}\{u_{k,n}\}$ at $\varepsilon=\varepsilon_{k,n}$, and
  3. $(iii)$
    the transversality condition $\mathcal{P}(\mathrm{d}_{1}\mathrm{d}_{2}\mathscr{F}[\varepsilon_{k,n},0](1,u_{k,n}))\neq 0$ holds, where $\mathcal{P}:L^{2}_{\mathrm{e}}\to L^{2}_{\mathrm{e}}$ is a projection with $\mathrm{Im}(\mathcal{L})=\mathrm{ker}(\mathcal{P}).$
Conditions $(i)$ and $(ii)$ were considered in Section 2 and so it remains to prove that condition $(iii)$ holds. We begin by noting the following characterisation of the projection operator $\mathcal{P}$.
Remark 3.1.
Suppose $\mathcal{L}$ is a symmetric operator with respect to some inner product $\langle\cdot,\cdot\rangle$; i.e,
$\langle\mathcal{L}u_{1},u_{2}\rangle=\langle u_{1},\mathcal{L}u_{2}\rangle,\qquad\qquad\forall u_{1},u_{2}\in X,$
and $\mathrm{ker}(\mathcal{L})=\mathrm{span}(v_{0})$. There exists a projection $\mathcal{P}$ with $\mathrm{Im}(\mathcal{P})=\mathrm{ker}(\mathcal{L})$ given by
$\mathcal{P}\,u=\frac{\langle v_{0},u\rangle}{\langle v_{0},v_{0}\rangle}\,v_{0}.$
We are now able to prove the following lemma regarding the transversality condition $(iii)$.
Lemma 3.2.
Suppose that $\mathrm{ker}\,(\mathcal{L})=\mathrm{span}\{v_{0}\}$ for some $\varepsilon_{0}>0$ and $v_{0}\in H^{4}_{\mathrm{e}}$. It follows that
$\mathcal{P}(\mathrm{d}_{1}\mathrm{d}_{2}\mathscr{F}[\varepsilon_{0},0](1,v_{0}))=0\qquad\iff\qquad v_{0}=0.$
Proof.
We prove this statement in two steps. We first show that $\mathcal{P}(\mathrm{d}_{1}\mathrm{d}_{2}\mathscr{F}[\varepsilon_{0},0](1,v_{0}))=0$ is equivalent to $(1+\Delta)v_{0}=0$ for $v_{0}\in\mathrm{ker}\,(\mathcal{L})$, before then proving that $(1+\Delta)v_{0}=0$ if and only if $v_{0}=0$ for any $v_{0}\in H^{4}_{\mathrm{e}}$.
We note that $\mathcal{L}$ is symmetric with respect to the standard real-valued $L^{2}$ inner product; i.e.,
$\langle\mathcal{L}u_{1},u_{2}\rangle_{L^{2}}=\langle u_{1},\mathcal{L}u_{2}\rangle_{L^{2}},\qquad\qquad\forall u_{1},u_{2}\in H^{4}_{\mathrm{e}}$
and so it follows that we can define $\mathcal{P}$ as
$\mathcal{P}\,u=\frac{\langle v_{0},u\rangle_{L^{2}}}{\|v_{0}\|^{2}_{L^{2}}}\,v_{0},$
as stated in Remark 3.1. Noting that $\mathrm{d}_{1}\mathrm{d}_{2}\mathscr{F}[\varepsilon_{0},0](1,v)=-2\varepsilon_{0}\,[-1+2\,V(|\mathbf{x}|)]\,v=\frac{2}{\varepsilon_{0}}(\mathcal{L}+(1+\Delta)^{2})v$, we obtain
[…]
and so $\mathcal{P}\left(\mathrm{d}_{1}\mathrm{d}_{2}\mathscr{F}[\varepsilon_{0},0](1,v_{0})\right)=0$ if and only if $v_{0}\,(1+\Delta)v_{0}=0$. For any function $v\in H^{1}(\mathbb{R}^{n};\mathbb{R})$ satisfying $(1+\Delta)v=0$, we note that Pokhozhaev’s identity (cf. Berestycki and Lions [6, (2.1)])
$\frac{n-2}{2}\|\nabla v\|_{L^{2}(\mathbb{R}^{n})}^{2}=\frac{n}{2}\|v\|_{L^{2}(\mathbb{R}^{n})}^{2}$
holds for each $n\in\mathbb{N}$ and thus, for $v_{0}\in H^{4}_{\mathrm{e}}$ satisfying $(1+\Delta)v_{0}=0$ we obtain
$\|v_{0}\|_{L^{2}}^{2}=0$
which implies that $v_{0}=0$. ∎
We now present our existence theorem for the local bifurcation curve.
Theorem 3.3 (Local bifurcation).
Fix $R>0$, $k\in\mathbb{N}_{0}$ and $n\in\mathbb{N}$ such that Hypothesis 1 holds for $\varepsilon=\varepsilon_{k,n}$. A bifurcation occurs at $\varepsilon=\varepsilon_{k,n}$; that is, for some $\delta>0$, there exist neighbourhoods $\mathcal{N}\subset\mathbb{R}$ of $\varepsilon_{k,n}$ and $\mathcal{V}\subset H^{4}_{\mathrm{e}}$ of 0, and analytic functions $\mathcal{E}:(-\delta,\delta)\to\mathcal{N}$, $v:(-\delta,\delta)\to\mathcal{V}$ such that all nontrivial solutions to (3.1) in $\mathcal{U}:=\mathcal{N}\times\mathcal{V}$ lie on the solution branch
$\mathcal{C}_{\delta}:=\left\{(\mathcal{E}(s),v(s)):s\in(-\delta,\delta)\right\},$ (3.2)
with $\mathcal{E}(0)=\varepsilon_{k,n}$, $v(0)=0$ and $v^{\prime}(0)=u_{k,n}$.
Proof.
The existence of a (unique) local bifurcation curve follows directly from Theorem A.1 with $X=H^{4}_{\mathrm{e}}$, $Y=L^{2}_{\mathrm{e}}$, $\varepsilon_{0}=\varepsilon_{k,n}$, $v_{0}=u_{k,n}$, and $\mathscr{F}(\varepsilon,u)$ given by (3.1). Condition $(i)$ follows from Lemma 2.1, $(ii)$ from Lemma 2.3 and Hypothesis 1, and $(iii)$ from Lemma 3.2. ∎
Having proved the local existence of bifurcating solutions to (1.2), we now present additional results regarding the characterisation of the local solution branch $\mathcal{C}_{\delta}$. We begin by considering the rotational invariance of solutions along each local bifurcation branch; we recall the rotational subgroups $X_{k}^{4}$ defined in (1.4), referring to Remark 1.1 for the cases $k=0,1$, and prove the following result.
Corollary 3.4.
Fix $R>0$, $k\in\mathbb{N}_{0}$ and $n\in\mathbb{N}$ as in Theorem $3.$ There exists some $0<\delta^{*}\leq\delta$ such that any solution $(\varepsilon,u)$ to (3.1) along the local bifurcation curve $\mathcal{C}_{\delta^{*}}\subseteq\mathcal{C}_{\delta}$ lies in the rotational subspace $\mathbb{R}\times X^{4}_{k}$.
Proof.
The existence of a local bifurcation curve $\mathcal{C}_{\delta^{*}}\subseteq\mathbb{R}\times X^{4}_{k}$ follows from respectively replacing $H^{4}_{\mathrm{e}}$ and $L^{2}_{\mathrm{e}}$ with $X^{4}_{k}$ and $X^{0}_{k}$ in the proof of Theorem $3.$ Since $X^{4}_{k}$ is a subspace of $H^{4}_{\mathrm{e}}$, we conclude that this bifurcation curve and the curve obtained in Theorem 3.3 are equivalent, by uniqueness in $H^{4}_{\mathrm{e}}$. ∎
Utilising the analytic properties of the local bifurcation curve, we also apply a supplementary theorem presented in Groves and Horn [30, Theorem 4.1] to characterise the local bifurcation curve $\mathcal{C}_{\delta}$. We state this result as Theorem A.2 in the appendix, which we apply in the following corollary.
Corollary 3.5.
Fix $R>0$, $k\in\mathbb{N}_{0}$ and $n\in\mathbb{N}$ as in Theorem 3.3, so that a local bifurcation occurs at $\varepsilon=\varepsilon_{k,n}$, with bifurcating profile $u(\mathbf{x})=s\,u_{k,n}(\mathbf{x})+\mathcal{O}(s^{2})$ for $|s|<\delta$. The local bifurcation is a pitchfork for all $k\neq 0$; for $k=0$ the bifurcation is a pitchfork if $f^{\prime\prime}(0)\,\langle u^{\vphantom{2}}_{k,n},u_{k,n}^{2}\rangle_{L^{2}}=0$ and a transcritical bifurcation otherwise.
Proof.
The characterisation of the local bifurcation curve follows directly from Theorem A.2, where we note that
$\varepsilon_{1}=-\frac{\mathcal{P}\left(\frac{1}{2!}\,\mathrm{d}_{2}^{2}\mathscr{F}[\varepsilon_{k,n},0](u_{k,n},u_{k,n})\right)}{\mathcal{P}\left(\mathrm{d}_{1}\mathrm{d}_{2}\mathscr{F}[\varepsilon_{k,n},0](1,u_{k,n})\right)}=-\frac{\varepsilon_{k,n}}{2}\,f^{\prime\prime}(0)\,\frac{\langle u^{\vphantom{2}}_{k,n},u_{k,n}^{2}\rangle_{L^{2}}}{\|(1+\Delta)u_{k,n}\|^{2}_{L^{2}}}.$
In particular, for $k\neq 0$ we obtain $u_{k,n}^{2}(r\cos(\theta),r\sin(\theta))=\frac{1}{2}v_{k,n}^{2}(r)+\frac{1}{2}v_{k,n}^{2}(r)\,\cos(2k\theta)$ and so, by the orthogonality of the angular Fourier modes, it follows that
$\langle u^{\vphantom{2}}_{k,n},u_{k,n}^{2}\rangle_{L^{2}}=0$
for any $k,n\in\mathbb{N}$. ∎
We now consider the stability of solutions along the local branch $\mathcal{C}_{\delta}$. In particular, we prove the spectral (in)stability of local bifurcating solutions for transcritical, subcritical, or supercritical bifurcations.
Lemma 3.6.
Fix $R>0$, $k\in\mathbb{N}_{0}$ and $n\in\mathbb{N}$ as in Theorem 3.3, so that a local bifurcation occurs at $\varepsilon=\varepsilon_{k,n}$, with bifurcating profile $u(\mathbf{x})=s\,u_{k,n}(\mathbf{x})+\mathcal{O}(s^{2})$ for $|s|<\delta$. If $\varepsilon=\varepsilon_{k,n}$ is the primary bifurcation point, then the bifurcation curve consists of
  1. (i)
    1 stable and 1 unstable branch if the bifurcation is transcritical;
  2. (ii)
    2 stable branches if the bifurcation is a supercritical pitchfork; and
  3. (iii)
    2 unstable branches if the bifurcation is a subcritical pitchfork.
Each subsequent local bifurcation curve consists of solely unstable branches.
Proof.
We only consider the primary bifurcation branch where the trivial state first destabilises; the eigenvalues of the bifurcating solution will be a small perturbation of the eigenvalues of the trivial state (cf. Kato [47, Theorem VIII-2.6]), and so each bifurcating solution will be unstable for an unstable trivial state.
We consider the eigenvalue problem for the linearisation of (1.2) about the solution $u=v(s)$
$\lambda w=\mathcal{L}_{v(s)}\,w=-(1+\Delta)^{2}w-\varepsilon^{2}w+2\varepsilon^{2}\,V(|\mathbf{x}|)\,w+f^{\prime}(v(s))\,w$ (3.3)
where we take $\lambda=\lambda^{*}$ to be the most unstable eigenvalue; i.e.,
$\mathrm{Re}(\lambda^{*})>\mathrm{Re}(\lambda)\qquad\text{for all \;$\lambda\in\sigma(\mathcal{L}_{v(s)})\backslash\{\lambda^{*}\}$}.$
Since we have assumed that the trivial state destabilises at the bifurcation point, it follows that $\lambda^{*}|_{s=0}=0$. We note that $\lambda^{*}|_{s=0}=0$ is simple, by Hypothesis 1, and so there exists an analytic function $\Lambda:(-\delta,\delta)\to\mathbb{R}$ such that $\lambda^{*}=s\,\Lambda(s)$ in a neighbourhood of the curve $\mathcal{C}_{\delta}$ (see Kato [47, VII-§3.2]). We consider the following Taylor expansions
$v(s)=u_{k,n}\,s+\sum_{n=1}^{\infty}v_{n}\,s^{n+1},\qquad\mathcal{E}(s)=\varepsilon_{k,n}+\sum_{n=1}^{\infty}\varepsilon_{n}\,s^{n},\qquad\Lambda(s)=\lambda_{0}+\sum_{n=1}^{\infty}\lambda_{n}\,s^{n},$
where $v_{1},v_{2},\dots\in\mathrm{ker}(\mathcal{Q})$, $\varepsilon_{1},\varepsilon_{2},\dots\in\mathbb{R}$ and $\lambda_{0},\lambda_{1},\dots\in\mathbb{R}$, such that
$u=v(s),\qquad\varepsilon=\mathcal{E}(s),\qquad\lambda=s\,\Lambda(s),\qquad s\in(-\delta,\delta)$
for some $\delta>0$. Expanding (3.3) in powers of $s\in(-\delta,\delta)$, we obtain
$\mathcal{O}(1):$ $\hskip 18.49988pt0={}$ $-(1+\Delta)^{2}w-\varepsilon_{k,n}^{2}w+2\,\varepsilon_{k,n}^{2}\,V(|\mathbf{x}|)\,w,$
$\mathcal{O}(s):$ $\hskip 18.49988pt\lambda_{0}w={}$ $-2\,\varepsilon_{k,n}\,\varepsilon_{1}w+4\,\varepsilon_{k,n}\,\varepsilon_{1}\,V(|\mathbf{x}|)\,w+f^{\prime\prime}(0)\,u_{k,n}\,w,$
$\mathcal{O}(s^{2}):$ $\lambda_{1}w={}$ $-\left(\varepsilon_{1}^{2}+2\varepsilon_{k,n}\,\varepsilon_{2}\right)\,[-1+2V(|\mathbf{x}|)]\,w+f^{\prime\prime}(0)\,v_{1}\,w+\frac{1}{2!}\,f^{\prime\prime\prime}(0)\,u_{k,n}^{2}\,w,$
The first equation yields $w=u_{k,n}$, resulting in the condition
$\lambda_{0}\langle u_{k,n},u_{k,n}\rangle_{L^{2}}={}$ $\varepsilon_{1}\,\langle u_{k,n},\mathrm{d}_{1}\mathrm{d}_{2}\mathscr{F}[\varepsilon_{k,n},0](1,u_{k,n})\rangle_{L^{2}}+f^{\prime\prime}(0)\,\langle u_{k,n},u_{k,n}^{2}\rangle_{L^{2}},$
and so
$\lambda_{0}={}$ $\frac{2\,\varepsilon_{1}}{\varepsilon_{k,n}}\,\frac{\|(1+\Delta)\,u_{k,n}\|^{2}_{L^{2}}}{\|u_{k,n}\|^{2}_{L^{2}}}+f^{\prime\prime}(0)\,\frac{\langle u_{k,n},u_{k,n}^{2}\rangle_{L^{2}}}{\|u_{k,n}\|^{2}_{L^{2}}}.$
Note the critical case when $\lambda_{0}=0$ is the condition that characterises a transcritical bifurcation in Theorem A.2. For $\lambda_{0}\neq 0$, we see that $\lambda=s\,\Lambda(s)$ changes sign at $s=0$ for $|s|\ll 1$, and so one bifurcating branch is stable and the other is unstable.
If $\langle u_{k,n},u_{k,n}^{2}\rangle_{L^{2}}=0$, we instead set $\varepsilon_{1}=0$ and obtain the leading-order condition
$\lambda_{1}={}$ $\frac{2\,\varepsilon_{2}}{\varepsilon_{k,n}}\frac{\|(1+\Delta)\,u_{k,n}\|_{L^{2}}^{2}}{\|u_{k,n}\|_{L^{2}}^{2}}+f^{\prime\prime}(0)\,\frac{\langle u_{k,n},u_{k,n}\,v_{1}\rangle_{L^{2}}}{\|u_{k,n}\|_{L^{2}}^{2}}+\frac{1}{2!}\,f^{\prime\prime\prime}(0)\,\frac{\|u_{k,n}\|_{L^{4}}^{4}}{\|u_{k,n}\|_{L^{2}}^{2}}$
where the critical case $\lambda_{1}=0$ is the condition that characterises a pitchfork bifurcation in Theorem A.2. If $f^{\prime\prime}(0)\,\langle u_{k,n},u_{k,n}\,v_{1}\rangle_{L^{2}}+\frac{1}{2!}\,f^{\prime\prime\prime}(0)\,\|u_{k,n}\|_{L^{4}}^{4}>0$, then the pitchfork bifurcation is subcritical and the bifurcating branches are unstable. Conversely, if $f^{\prime\prime}(0)\,\langle u_{k,n},u_{k,n}\,v_{1}\rangle_{L^{2}}+\frac{1}{2!}\,f^{\prime\prime\prime}(0)\,\|u_{k,n}\|_{L^{4}}^{4}<0$, then the pitchfork bifurcation is supercritical and the bifurcating branches are stable. ∎
Figure 6: Stability of the primary bifurcating branch for (a) transcritical, (b) supercritical pitchfork, and (c) subcritical pitchfork bifurcations from the trivial state. Here solid and dashed lines indicate stable and unstable solutions, respectively.
Thus, we obtain the expected stability behaviour associated with transcritical and pitchfork bifurcations; the axisymmetric solution bifurcates along a stable and unstable branch when $f^{\prime\prime}(0)\langle u^{\vphantom{2}}_{k,n},u_{k,n}^{2}\rangle_{L^{2}}\neq 0$, whereas the non-axisymmetric branch (or the axisymmetric branch with $f^{\prime\prime}(0)\langle u^{\vphantom{2}}_{k,n},u_{k,n}^{2}\rangle_{L^{2}}=0$) bifurcates along two stable or two unstable branches, depending on the criticality of the pitchfork; see Figure 6.
Remark 3.7.
We note that Lemma 3.6 holds when restricted to any rotational subspace $X^{4}_{k}$ that $\mathcal{C}_{\delta}$ lies in. This provides additional information regarding the stability of bifurcating solutions with respect to different dihedral perturbations.
As an example, suppose the first two bifurcation points are of the form $\varepsilon=\varepsilon_{1,1},\;\varepsilon_{0,1}$, with each respective solution $u=u_{1,1},\;u_{0,1}$ bifurcating supercritically. Then, in the full space $H^{4}_{\mathrm{e}}$ Lemma 3.6 states that the solution $u=u_{1,1}$ is linearly stable close to its bifurcation point and $u=u_{0,1}$ is unstable. If we instead restrict to the space of axisymmetric functions, then $u=u_{1,1}$ is no longer a valid solution and so $u=u_{0,1}$ is the primary branch and thus linearly stable by Lemma 3.6. It follows that, in this case, $u=u_{0,1}$ is linearly stable with respect to all dihedral and axisymmetric perturbations (i.e. perturbations within the $\mathbb{D}_{k}$ symmetry group for $k\neq 1$) but is unstable in $H^{4}_{\mathrm{e}}$, while the solution $u=u_{1,1}$ is linearly stable for all perturbations in $H^{4}_{\mathrm{e}}$.
Remark 3.8.
For particular choices of the nonlinearity $f(u)$, we can immediately predict the bifurcation structure of these localised patterns. In particular, for $f(u)=-u^{3}$ (or, indeed, any odd analytic function $f$ with $f^{\prime\prime\prime}(0)<0$) every bifurcation will be a supercritical pitchfork, and so the primary bifurcating branches will both be stable.
We augment our analysis with numerical simulations (using codes available at [38]) of (1.2) with $f(u)=\nu u^{2}-u^{3}$, which we present in Figure 7. We consider three cases $R=2.6,2.8,7.5$ for the width of the potential $V$, corresponding to the three regimes highlighted in Figure 5, and two choices $\nu=0,0.4$ of the quadratic coefficient. As predicted, we observe that the primary bifurcation branch is linearly stable to the right of the bifurcation point and unstable to the left, and Figures 7(a-b) and (d-e) show that the primary branch changes from the $\mathbb{D}_{0}$ spot to the $\mathbb{D}_{1}$ dipole pattern as $R$ passes from $2.6$ to $2.8$. Furthermore, Figures 7(c) and (f) show that a $\mathbb{D}_{6}$ branch bifurcates for $R=7.5$, as predicted in Figure 5(d).
We note that, by including the quadratic coefficient $\nu=0.4$, the $\mathbb{D}_{0}$ branch changes from a supercritical pitchfork to a transcritical bifurcation, while the $\mathbb{D}_{1}$ branch changes from a supercritical to a subcritical pitchfork when $R=2.8$; in contrast, the $\mathbb{D}_{1}$ branch remains unchanged for $R=2.6$, as does the $\mathbb{D}_{6}$ branch for $R=7.5$. While the $\mathbb{D}_{1}$ branch is qualitatively the same in Figures 7(a) and (d), and likewise the $\mathbb{D}_{6}$ branch in Figures 7(c) and (f), we emphasise that the solution profiles qualitatively change as $\nu$ increase from 0 to $0.4$. We note that when $R=2.8$, such that the primary branch is the $\mathbb{D}_{1}$ dipole pattern, the $\mathbb{D}_{0}$ branch is initially unstable but then restabilises as a pair of unstable $\mathbb{D}_{1}$ branches emerge from a secondary bifurcation; see Figures 7(b) and (e).
Figure 7: Numerical simulations of (1.2) with $f(u)=\nu\,u^{2}-u^{3}$. Columns indicate (Left) $R=2.6$, (Centre) $R=2.8$, and (Right) $R=7.5$, while rows indicate (Top) $\nu=0$ and (Bottom) $\nu=0.4$. The blue, red and green curves are the solution branches for localised $\mathbb{D}_{0}$, $\mathbb{D}_{1}$ and $\mathbb{D}_{6}$ patterns, respectively, and solid/dashed lines indicate linearly stable/unstable solutions.
3.2 Global bifurcation
We now prove that the local bifurcation curves found in the previous section can be continued away from the bifurcation point; we call these global bifurcation branches, and use the analytic global implicit function theorem of Chen, Walsh and Wheeler [17, Theorem B.1] (which we present as Theorem A.3 in the appendix) in order to prove their existence.
Theorem 3.9 (Global bifurcation).
Fix $R>0$, $k\in\mathbb{N}_{0}$ and $n\in\mathbb{N}$ as in Theorem 3.3, so that there exists a local bifurcation curve $\mathcal{C}_{\delta}$ passing through the point $(\varepsilon,u)=(\varepsilon_{k,n},0)$. Then, there exists an open set $U\subset\mathbb{R}\times H^{4}_{\mathrm{e}}$, with $(\varepsilon_{k,n},0)\in U$, such that $\mathcal{C}_{\delta}$ can be extended to a global bifurcation curve $\mathcal{C}=\{(\mathcal{E}(s),v(s))\,:\;s\in\mathbb{R}\}$ of solutions to (3.1), where $(\mathcal{E},v):\mathbb{R}\to U$ is a continuous function. Furthermore,
  1. (a)
    $\mathcal{C}$ has a local analytic parametrisation around each of its points;
  2. (b)
    any locally analytic bifurcation curve in $U$ which includes the point $(\varepsilon_{k,n},0)$ is contained in $\mathcal{C}$; and
  3. (c)
    one of the following alternatives occurs:
    1. I:
      $\mathcal{C}$ is unbounded;
    2. II:
      $\mathcal{C}$ approaches the boundary of $U$;
    3. III:
      $\mathcal{C}$ forms a closed loop; or
    4. IV:
      $\mathcal{C}$ loses compactness.
Proof.
This result follows directly from Theorem A.3 with $X=H^{4}_{\mathrm{e}}$ and $Y=L^{2}_{\mathrm{e}}$. While the bifurcation point $(\varepsilon_{k,n},0)$ does not satisfy the conditions for Theorem A.3, since $\mathrm{d}_{2}\mathscr{F}[\varepsilon_{k,n},0]$ is not an isomorphism $H^{4}_{\mathrm{e}}\to L^{2}_{\mathrm{e}}$, it follows that there is a point $(\varepsilon_{*},v_{*})\in\mathcal{C}_{\delta}$ for which Theorem A.3 holds. This is due to the Fredholm properties of $\mathcal{L}_{\phi}:=\mathrm{d}_{2}\mathscr{F}[\varepsilon,\phi]$ proved in Lemma 2.1 for any $(\varepsilon,\phi)\in U$, as well as the transversality property proved in Lemma 3.2. The possibility of a loss of Fredholm properties in Theorem A.3 is also avoided since we proved that $\mathrm{d}_{2}\mathscr{F}[\varepsilon,\phi]$ is Fredholm for all $\varepsilon>0$, $\phi\in H^{4}_{\mathrm{e}}$ in Lemma 2.1. ∎
We make the following remarks regarding Theorem 3.9.
Remark 3.10.
  1. (i)
    Much like in Corollary 3.4, Theorem 3.9 holds when replacing the respective spaces $H^{4}_{\mathrm{e}}$ and $L^{2}_{\mathrm{e}}$ with $X^{4}_{k}$ and $X^{0}_{k}$, and so the global bifurcation curve is also contained in the rotational subspace $\mathbb{R}\times X^{4}_{k}$.
  2. (ii)
    We note that $\mathcal{C}$ having an analytic parametrisation at each point does not imply that $\mathcal{C}$ is smooth. In fact, as discussed in [11, Remark 9.1.2 (3)], the curve $\mathcal{C}$ can possess isolated points where $\mathcal{C}$ is not even continuously differentiable. We demonstrate an example where $\mathcal{C}$ has an isolated corner in Figure 8(iv).
  3. (iii)
    Different versions of the alternatives I-III are presented in the literature (see, for example [13, Remarks 4.2 and 4.3]); in particular, Constantin, Strauss and Vărvărucă [18, Remark 8] highlight that the alternatives presented by Buffoni and Toland [11, Theorem 9.1.1] are not exhaustive, and present a different formulation in their global bifurcation theorem [18, Theorem 6]. Here we use the same formulation of the alternatives as used by Chen, Walsh and Wheeler [17, Theorem B.1], which we present in more detail in Theorem A.3. Note that alternatives $I$ and $II$ are often indistinguishable from one another and so we combine the two into alternative $(i)$ in Theorem A.3.
  4. (iv)
    In many applications one can exclude the case when $\mathcal{C}$ forms a closed loop in part $(c)$, typically via bifurcation theory on cones [11, Theorem 9.2.2] and using nodal properties of the bifurcating solution. However, in our problem we numerically observe cases where $\mathcal{C}$ forms a closed loop (such as in Figure 8(iv)) and so we do not expect this to be possible in general.
  5. (v)
    The loss of compactness alternative can sometimes be excluded through certain properties of solutions; for example, Haziot and Wheeler [32] used the monotonicity of their solutions while Kozlov [49] utilised uniform decay rates. We have been unable to exclude a loss of compactness in the current setting, but it may be possible to further strengthen our global bifurcation result in future studies.
We again provide numerical simulations to support our analysis, which we present in Figure 8. For fixed $R=2.8$, we can vary $\nu$ such that we obtain the three different alternatives in Theorem 3.9 for the primary $\mathbb{D}_{1}$ bifurcation. As such, we are unable to disregard any of the possible alternatives presented in Theorem 3.9.
For $\nu\neq 0$, the dipole pattern evolves into a combination of a radial spot and the dipole; for $\nu=0.4$ this results in a loop connecting the bifurcation point to a secondary bifurcation at the radial spot, while for $\nu=1.6$ the branch converges towards $\varepsilon=0$. In this case the solution returns to a dipole solution but with weakening localisation, as plotted in Figure 8 $(vi)$. We note that each branch plotted in Figure 8 is actually two branches, the solution and its half-period rotation, which is why the closed loop appears to be a single branch terminating at a point. We also note that the point $(iv)$ in Figure 8 forms an isolated corner on the closed loop, which is permitted within Theorem 3.9 (see Remark 3.10(iii)(iii)).
Figure 8: Numerical simulation of the primary branch, consisting of a $\mathbb{D}_{1}$ dipole pattern, for (1.2) with $f(u)=\nu u^{2}-u^{3}$, $R=2.8$, and $\nu=0,0.4,1.6$. Line style and colour of branches indicates the same properties as in Figure 7. The solution $(i\to ii)$ blows up in its norm for $\nu=0$, $(i\to iii\to iv)$ forms a closed loop for $\nu=0.4$, and $(i\to v\to vi)$ approaches the boundary $\varepsilon=0$ for $\nu=1.6$.
4 Discussion
In this paper we have proved the existence of large-amplitude, fully localised 2D patterns in the Swift–Hohenberg equation with an axisymmetric spatial heterogeneity. It is somewhat remarkable that one is able to do this despite there being no equivalent results in the spatially homogeneous case; indeed, the spatial heterogeneity considered in this paper allows us to derive almost explicit characterisations of the shape of the 2D localised pattern and its linear stability near bifurcation. Despite choosing an axisymmetric heterogeneity, non-axisymmetric localised solutions can emerge and even bifurcate as a stable localised pattern. This raises the question of whether such patterns can be realised experimentally.
This work lays the foundation for further exploration of localised patterns induced by compact spatial heterogeneities. A straightforward extension would be to consider the three-dimensional analogue of (1.2), where the Fredholm properties will remain the same and the point spectrum can again be computed explicitly using spherical harmonics and spherical Bessel functions. Replacing the step function $V$ by a smooth potential would be a more realistic model for spatial heterogeneities in experiments, and one would expect much of the underlying theory of this work to persist in that case. In particular, we highlight the approach presented by Brooks, Derks and Lloyd [9, Section 5] in which they prove the persistence of fronts in the sine-Gordon equation from piecewise to smooth heterogeneities; we expect a similar argument to hold in this work using the approach of van Heijster and Sandstede [62].
Different types of spatial heterogeneity can be considered, for instance, non-axisymmetric heterogeneities like those studied in [41, 40] or an axisymmetric heterogeneity translating uniformly in space. The most challenging part in these cases is then to explicitly compute the point spectrum of the linear operator. Another common type of spatial heterogeneity is when $u$ is no longer multiplying $V$, e.g.
$\partial_{t}u=-(1+\Delta)^{2}u-\varepsilon^{2}u+f(u)+\delta\,V(\mathbf{x}).$
This type of heterogeneity is more similar to the form of heterogeneity found in the Rayleigh–Bénard convection problem, where a non-uniform heat profile is applied to the bottom of the fluid. This problem has the additional difficulty that the trivial state is no longer an equilibrium state, but the small defect case (i.e. with $|\delta|\ll 1$) may still be amenable to methods presented here.
We recall that our approach is inspired by experimental and numerical studies, where a spatial heterogeneity is added to the system in order to induce the emergence of localised patterns. However, there is a key element of this approach which we have not considered here, namely, the persistence of these patterns. In the ferrofluid experiment, for example, the spatial heterogeneity is subsequently removed from the system and the induced localised spikes remain on the fluid surface. This represents a method by which one can find localised patterns in a spatially homogeneous system by adding and removing a compact heterogeneity, which we would like to replicate in our analysis. Mathematically, this is equivalent to finding $(\mathcal{E}(s),v(s))\in\mathcal{C}$ on the global bifurcation curve for (1.2), and then seeking solutions to the homogeneous Swift–Hohenberg equation (i.e. (1.2) with $V\equiv 0$) of the form $\varepsilon=\mathcal{E}(s)$, $u=v(s)+w$ with $w\in H^{4}_{\mathrm{e}}$. In particular, one would need to solve
$(1+\Delta)^{2}+\mathcal{E}^{2}-f^{\prime}(v)]w=-2\,\mathcal{E}^{2}\,V\,v+N(w;v)$
for $w\in H^{4}_{\mathrm{e}}$, where $N(w;v):=f(v+w)-f(v)-f^{\prime}(v)\,w=\mathcal{O}(|w|^{2})$, possibly via a Newton–Kantorovich or implicit function argument. Numerically, we observe that the remainder function $w$ is not small relative to $v$, making it difficult to rule out the trivial solution $w\equiv-v$ (corresponding to $u\equiv 0$). It may still be possible to prove the existence of a non-trivial solution—thereby solving the open problem of proving the existence of fully localised 2D patterns—but we leave this as a future avenue to explore.
We expect the approach presented in this work to be easily extended to other types of PDE systems, and so it would be natural to consider models that are more specific to particular experiments or phenomena. As discussed, the ferrofluid experiment studied by Richter and Barashenkov [59] is a key motivation for this work, and so it would be interesting to try and apply this theory to the ferrofluid problem. We note that there have already been studies of two-dimensional neural field models with spatial heterogeneity [58], including the case where firing rate is piecewise smooth [3, 8, 55], and so this may be a useful application to explore in greater detail.
Data availability statement
The data that support the findings of this study are openly available at [38].
Acknowledgments
The authors are grateful to Björn de Rijk, Mark D. Groves and Miles Wheeler for their assistance on certain aspects of this manuscript. DJH acknowledges support from the Alexander von Humboldt Foundation, and MT and DJBL acknowledge support from EPRSC grant UKRI070. The authors also thank Tom Bridges, Ryan Goh, Bastian Hilder and Jörg Weber for helpful discussions and suggestions regarding this work. For the purpose of Open Access, the authors have applied a Creative Commons Attribution (CC BY) public copyright licence to any Author Accepted Manuscript version arising from this submission.
Appendix A Bifurcation theorems
There are several bifurcation theorems that we have used in this work, each of which is a standard tool in bifurcation theory. To balance accessibility and brevity, we present the full statement of each theorem that we use and provide further references for those interested in their proofs.
In our local bifurcation theorem (Theorem 3.3) we use an analytic version of the Crandall–Rabinowitz theorem (cf. Buffoni and Toland [11, Theorem 8.3.1]), which we state here with the notation presented by Groves and Horn [30, Theorem 3.1].
Theorem A.1 (Crandall–Rabinowitz theorem).
Let $X$ and $Y$ be Banach spaces, $V$ be an open neighbourhood of the origin in $X$ and $\mathscr{F}:\mathbb{R}\times V\to Y$ be an analytic function with $\mathscr{F}(\varepsilon,0)=0$ for all $\varepsilon\in\mathbb{R}$. Suppose also that
  1. (i)
    $\mathcal{L}:=\mathrm{d}_{2}\mathscr{F}[\varepsilon_{0},0]:X\to Y$ is a Fredholm operator of index zero,
  2. (ii)
    $\mathrm{ker}(\mathcal{L})=\mathrm{span}\{v_{0}\}$ for some $v_{0}\in X$, and
  3. (iii)
    the transversality condition $\mathcal{P}(\mathrm{d}_{1}\mathrm{d}_{2}\mathscr{F}[\varepsilon_{0},0](1,v_{0}))\neq 0$ holds, where $\mathcal{P}:Y\to Y$ is a projection with $\mathrm{Im}(\mathcal{L})=\mathrm{ker}(\mathcal{P}).$
The point $(\varepsilon_{0},0)$ is a local bifurcation point, that is there exist $\delta>0$, an open neighbourhood $W$ of $(\varepsilon_{0},0)$ in $\mathbb{R}\times X$ and analytic functions $v:(-\delta,\delta)\to V$, $\mathcal{E}:(-\delta,\delta)\to\mathbb{R}$ with $\mathcal{E}(0)=\varepsilon_{0}$, $v(0)=0$, $v^{\prime}(0)=v_{0}$ such that $\mathscr{F}(\mathcal{E}(s),v(s))=0$ for every $s\in(-\delta,\delta)$. Furthermore,
$W\cap N=\{(\mathcal{E}(s),v(s)):0<|s|<\delta\},$
where
$N=\{(\varepsilon,u)\in\mathbb{R}\times V\backslash\{0\}:\mathscr{F}(\varepsilon,u)=0\}.$
A proof of the standard Crandall–Rabinowitz theorem can be found in the original paper by Crandall and Rabinowitz [19, Theorem 1], while the analytic version is proved by Buffoni and Toland [11, Theorem 8.3.1]. Using the analytic properties of the local bifurcation curve, we can also apply the following supplementary theorem presented by Groves and Horn [30, Theorem 4.1] to characterise the local bifurcation curve $\mathcal{C}_{\delta}$.
Theorem A.2.
Suppose that the hypotheses of Theorem A.1 hold. In the notation of that theorem, let $\mathcal{Q}:X\to X$ be a projection with $\mathrm{Im}(\mathcal{Q})=\mathrm{ker}(\mathcal{L})$ and the Taylor series of the functions $v:(-\delta,\delta)\to V$, $\mathcal{E}:(-\delta,\delta)\to\mathbb{R}$ be given by
$v(s)=v_{0}\,s+\sum_{n=1}^{\infty}v_{n}\,s^{n+1},\qquad\mathcal{E}(s)=\varepsilon_{0}+\sum_{n=1}^{\infty}\varepsilon_{n}\,s^{n},$
where $v_{1},v_{2},\dots\in\mathrm{ker}(\mathcal{Q})$ and $\varepsilon_{1},\varepsilon_{2},\dots\in\mathbb{R}$.
  1. (i)
    The coefficient $\varepsilon_{1}$ satisfies the equation
    $\mathcal{P}\left(\frac{1}{2!}\,\mathrm{d}_{2}^{2}\mathscr{F}[\varepsilon_{0},0](v_{0},v_{0})\right)+\varepsilon_{1}\mathcal{P}\left(\mathrm{d}_{1}\mathrm{d}_{2}\mathscr{F}[\varepsilon_{0},0](1,v_{0})\right)=0,$
    and the bifurcation is transcritical if $\varepsilon_{1}$ is non-zero.
  2. (ii)
    Suppose that $\varepsilon_{1}=0$. The coefficient $\varepsilon_{2}$ satisfies the equation
    $\mathcal{P}\left(\mathrm{d}_{2}^{2}\mathscr{F}[\varepsilon_{0},0](v_{0},v_{1})+\frac{1}{3!}\,\mathrm{d}_{2}^{3}\mathscr{F}[\varepsilon_{0},0](v_{0},v_{0},v_{0})\right)+\varepsilon_{2}\mathcal{P}\left(\mathrm{d}_{1}\mathrm{d}_{2}\mathscr{F}[\varepsilon_{0},0](1,v_{0})\right)=0,$
    where $v_{1}\in\mathrm{ker}(\mathcal{Q})$ solves the equation
    $\mathrm{d}_{2}\mathscr{F}[\varepsilon_{0},0](v_{1})=-\frac{1}{2!}\,\mathrm{d}_{2}^{2}\mathscr{F}[\varepsilon_{0},0](v_{0},v_{0})$
    The bifurcation is supercritical for $\varepsilon_{2}>0$ and subcritical for $\varepsilon_{2}<0$.
This theorem uses the analytic properties of bifurcating solutions in order to take Taylor series expansions for $|s|\ll 1$ and characterise the bifurcation structure by standard normal forms in bifurcation theory (see, for example Verhulst [63, Section 13.5]).
To prove our global bifurcation result (Theorem 3.9) we use the analytic global implicit function theorem given by Chen, Walsh and Wheeler [17, Theorem B.1], which extends the global bifurcation theorem of Buffoni and Toland [11, Theorem 9.1.1] to cases where compactness and Fredholm properties of solutions are no longer assumed.
Theorem A.3 (Global continuation).
Let $X$ and $Y$ be Banach spaces, $U\subseteq\mathbb{R}\times X$ an open set containing a point $(\varepsilon_{*},u_{*})$. Suppose that $\mathscr{F}:U\to Y$ is real analytic and satisfies
$\mathscr{F}(\varepsilon_{*},u_{*})=0,\qquad\qquad\mathrm{d}_{2}\mathscr{F}[\varepsilon_{*},u_{*}]\;\text{is an isomorphism $X\to Y$.}$
Then there exists a curve $\mathcal{C}$ that admits the global $C^{0}$ parametrisation
$\mathcal{C}:=\left\{(\mathcal{E}(s),v(s))\;:\;\;s\in\mathbb{R}\right\}\subseteq\mathscr{F}^{-1}(0)\cap U,$
and satisfies the following.
  1. (a)
    At each $s\in\mathbb{R}$, the linear operator $\mathrm{d}_{2}\mathscr{F}[\mathcal{E}(s),v(s)]:X\to Y$ is Fredholm with index zero.
  2. (b)
    One of the following alternatives holds as $s\to\infty$ and $s\to$-$\infty$.
    1. (i)
      (Blow up) The quantity
      $N(s):=\|v(s)\|_{X}+|\mathcal{E}(s)|+\frac{1}{\mathrm{dist}((\mathcal{E}(s),v(s)),\partial U)}\to\infty$
    2. (ii)
      (Loss of compactness) There exists a sequence $\{s_{n}\}$ with $s_{n}\to\pm\infty$ and $\sup_{n}N(s_{n})<\infty$, but $(\mathcal{E}(s_{n}),v(s_{n}))$ has no convergent subsequence in $\mathbb{R}\times X$.
    3. (iii)
      (Loss of Fredholmness) There exists a sequence $\{s_{n}\}$ with $s_{n}\to\pm\infty$, $\sup_{n}N(s_{n})<\infty$, and $(\mathcal{E}(s_{n}),v(s_{n}))\to(\mathcal{E}_{0},v_{0})\in U$ in $\mathbb{R}\times X$, however $\mathrm{d}_{2}\mathscr{F}[\mathcal{E}_{0},v_{0}]$ is not Fredholm with index zero.
    4. (iv)
      (Closed loop) There exists $T>0$ such that $(\mathcal{E}(s+T),v(s+T))=(\mathcal{E}(s),v(s))$ for all $s>0$.
  3. (c)
    Near each point $(\mathcal{E}(s_{0}),v(s_{0}))\in\mathcal{C}$, we can locally reparametrise $\mathcal{C}$ so that $s\mapsto(\mathcal{E}(s),v(s))$ is real analytic.
  4. (d)
    The curve $\mathcal{C}$ is maximal in the sense that, if $\mathcal{R}\subseteq\mathscr{F}^{-1}(0)\cap U$ is a locally real-analytic curve containing $(\varepsilon_{*},u_{*})$ and along which $\mathrm{d}_{2}\mathscr{F}$ is Fredholm of index zero, then $\mathcal{R}\subseteq\mathcal{C}$.
For a proof of this theorem, we refer the reader to Chen, Walsh and Wheeler [17, Theorem B.1], along with Buffoni and Toland [11, Theorem 9.1.1] for additional details.
Appendix B Additional proofs
We present several proofs regarding the spectral properties of the linear operator $\mathcal{L}_{\phi}$ given in (2.2).
We begin with a proof that multiplication by an element in $H^{m}=H^{m}(\mathbb{R}^{2})$ is a compact operator from $H^{m}$ to $L^{2}=L^{2}(\mathbb{R}^{2})$ with $m>1$; while this result is often used in linear stability analysis (cf. Kapitula and Promislow [46, Theorem 3.1.11]), we present a proof here for completeness.
Lemma B.1.
Suppose that $f\in H^{m}$ with $m>1$. The mapping $u\mapsto f\,u$ is a compact operator $H^{m}\to L^{2}$.
Proof.
Let $\{u_{n}\}$ be bounded in $H^{m}$, so that $\{f\,u_{n}\}$ is weakly convergent in $H^{m}$. Denote its weak limit by $g$ and note that $f\,u_{n}\to g$ (up to subsequences) in $L^{2}(B_{C})$ for any $C>0$, where $B_{C}$ denotes the unit ball of radius $C$ about the origin. Then, we obtain
[…]
for sufficiently large $n$. First choose $C$ sufficiently large such that the second and third terms are each less than $\frac{\varepsilon^{2}}{4}$, using the fact that $\|u_{n}\|_{\infty}$ is uniformly bounded since $H^{m}$ is continuously embedded in $L^{\infty}$. It follows that there exists some constant $N=N(C)>0$ such that the first term is less than $\frac{\varepsilon^{2}}{2}$ for all $n>N$, since $f\,u_{n}\to g$ in $L^{2}(B_{C})$. ∎
Additionally, we prove Lemma 2.3, characterising the eigenvalues and eigenfunctions of the linear eigenvalue problem (2.3). For the sake of readability, we first restate the lemma as originally presented in Section 2.2.
Lemma 2.3.
Fix $R>0$ and $\varepsilon>0$. The eigenvalue problem (2.3) possesses a solution $u=u_{k}\in H^{4}_{\mathrm{e}}$ given by $u_{k}(r\cos\theta,r\sin\theta)=v_{k}(r)\,\cos(k\theta)$ with
[…]
where $\alpha_{\pm}:=\sqrt{1\pm\varepsilon\,\sqrt{1-\tilde{\lambda}}}$, $\beta:=\sqrt{1+\mathrm{i}\,\varepsilon\,\sqrt{1+\tilde{\lambda}}}$; $J_{k},H_{k}^{(1)}$ are the $k$’th order Bessel functions of the first and third kind; $\psi$ is given by $\psi=\frac{1}{2}\sin^{-1}(\tilde{\lambda})+\frac{\pi}{4}$; $\phi$ is given by
$\mathrm{e}^{2\mathrm{i}\phi}=\mathrm{i}\,\mathrm{e}^{-\mathrm{i}\sin^{-1}(\tilde{\lambda})}\frac{W[J_{k}(\alpha_{+}r),\overline{H^{(1)}_{k}(\beta r)}](R)\,H^{(1)}_{k}(\beta R)}{W[J_{k}(\alpha_{+}r),H^{(1)}_{k}(\beta r)](R)\,\overline{H^{(1)}_{k}(\beta R)}};$
$W[\cdot,\cdot]$ denotes the weighted Wronskian function given by
$W[u,v](r):=r\left(u(r)\,v^{\prime}(r)-u^{\prime}(r)\,v(r)\right);$
and $\tilde{\lambda}=\tilde{\lambda}(R,\varepsilon)$ satisfies
$F_{k}(R,\varepsilon,\tilde{\lambda})=\mathrm{Re}\left[\mathrm{e}^{\mathrm{i}\sin^{-1}(\tilde{\lambda})}\,W[J_{k}(\alpha_{+}r),H^{(1)}_{k}(\beta r)](R)\,W[J_{k}(\alpha_{-}r),\overline{H^{(1)}_{k}(\beta r)}](R)\right]=0$
with $\tilde{\lambda}\notin(-\infty,-1]$.
Proof of Lemma 2.3.
We first decompose $u$ into an angular Fourier expansion
$u(r\cos\theta,r\sin\theta)=u_{0}(r)+\sum_{k=1}^{\infty}u_{k}(r)\,\cos(k\theta),$
with $u_{k}\in H^{4}_{(k)}((0,\infty))$ as defined by Groves and Hill [29] for each $k\in\mathbb{N}_{0}$. The linear problem (2.3) then becomes
$\tilde{\lambda}\,\varepsilon^{2}u_{k}=-(1+\mathcal{D}_{1-k}\mathcal{D}_{k})^{2}u_{k}-\varepsilon^{2}u_{k}+2\,\varepsilon^{2}\,V(r)\,u_{k},\qquad\qquad\forall\;k\in\mathbb{N}_{0},$
where we use the Bessel differential operators $\mathcal{D}_{k}:=\partial_{r}+\frac{k}{r}$ also introduced in [29]. For notational simplicity, we also define the $k$-index Laplace operator as $\Delta_{k}:=\mathcal{D}_{1-k}\mathcal{D}_{k}=\partial_{r}^{2}+\frac{1}{r}\partial_{r}-(\frac{k}{r})^{2}$. We seek solutions to (2.3) for an arbitrary $k\in\mathbb{N}_{0}$; to do this, we construct real bounded solutions on the disjoint regions $r<R$ and $r>R$ and match these functions at the discontinuity point $r=R$.
In both regions $r<R$ and $r>R$, the linear problem (2.3) can be written as
$0=-(1+\Delta_{k})^{2}u_{k}+\mu u_{k}$
for some $\mu\in\mathbb{R}$, which has a general solution of the form
$u_{k}=\tilde{A}_{1}J_{k}(\nu_{+}\,r)+\tilde{A}_{2}J_{k}(\nu_{-}\,r)+\tilde{A}_{3}Y_{k}(\nu_{+}\,r)+\tilde{A}_{4}Y_{k}(\nu_{-}\,r),\qquad\qquad\nu_{\pm}:=\sqrt{1\pm\sqrt{\mu}},$
where $J_{k},Y_{k}$ are $k$’th order Bessel functions of the first and second kind. Hence, any real bounded solution $u_{k}$ of (2.3) takes the form
[…]
with $A_{1},A_{2}\in\mathbb{R}$ and $A_{3}\in\mathbb{C}$, where $\alpha_{\pm}=\sqrt{1\pm\varepsilon\sqrt{1-\tilde{\lambda}}}$, $\beta=\sqrt{1+\mathrm{i}\varepsilon\sqrt{1+\tilde{\lambda}}}$, and $H^{(1)}_{k}$ is the $k$’th order Hankel function, given by
$H^{(1)}_{k}(x)=J_{k}(x)+\mathrm{i}Y_{k}(x).$
We decompose $\tilde{\lambda}$ and $\beta$ into their real and imaginary parts, which we denote as $\lambda_{R},\lambda_{I}$ and $\beta_{R},\beta_{I}$, respectively. The Hankel function $H_{k}^{(1)}$ then satisfies the following asymptotic behaviour
[…]
for $r\gg 1$ and, using the formula $\sqrt{z}=\sqrt{|z|}\,\frac{z+|z|}{|z+|z||}$ for any $z\in\mathbb{C}$, we note that
$\begin{split}\beta_{I}={}&\frac{\varepsilon\,\sqrt{(1+\lambda_{R})\,+\sqrt{(1+\lambda_{R})^{2}+\lambda_{I}^{2}}}}{2\,\sqrt{1-\frac{\varepsilon\,\lambda_{I}}{\sqrt{2}\,\sqrt{(1+\lambda_{R})\,+\sqrt{(1+\lambda_{R})^{2}+\lambda_{I}^{2}}}}+\sqrt{1+\varepsilon^{2}\,\sqrt{(1+\lambda_{R})^{2}+\lambda_{I}^{2}}-\frac{\sqrt{2}\,\varepsilon\,\lambda_{I}}{\sqrt{(1+\lambda_{R})\,+\sqrt{(1+\lambda_{R})^{2}+\lambda_{I}^{2}}}}}}}.\end{split}$
Hence, we observe that $\beta_{I}>0$ if and only if $1+\lambda_{R}+\sqrt{(1+\lambda_{R})^{2}+\lambda_{I}^{2}}>0$, and so $\tilde{\lambda}\notin(-\infty,-1]$ implies that bounded solutions $u_{k}$ to the linear problem (2.3) exhibit exponential radial decay and thus lie in $H^{4}_{(k)}(0,\infty)$. In order that a piecewise solution consisting of $u_{-}$ for $r<R$ and $u_{+}$ for $r>R$ is at least $C^{3}$ at $r=R$, the following matching conditions
$u_{-}(R)={}$ $u_{+}(R),$ $\hskip 18.49988pt\mathcal{D}_{k}u_{-}(R)={}$ $\mathcal{D}_{k}u_{+}(R),$
$\Delta_{k}u_{-}(R)={}$ $\Delta_{k}u_{+}(R),$ $\hskip 18.49988pt\mathcal{D}_{k}\Delta_{k}u_{-}(R)={}$ $\mathcal{D}_{k}\Delta_{k}u_{+}(R),$
must be satisfied. Hence, the constants $A_{1},A_{2},A_{3}$ must satisfy
$A_{1}u_{1}(R)+A_{2}u_{2}(R)={}$ $A_{3}u_{3}(R)+\overline{A_{3}}\,\overline{u_{3}}(R),$
$A_{1}\mathcal{D}_{k}u_{1}(R)+A_{2}\mathcal{D}_{k}u_{2}(R)={}$ $A_{3}\mathcal{D}_{k}u_{3}(R)+\overline{A_{3}}\,\mathcal{D}_{k}\overline{u_{3}}(R),$
$A_{1}\Delta_{k}u_{1}(R)+A_{2}\Delta_{k}u_{2}(R)={}$ $A_{3}\Delta_{k}u_{3}(R)+\overline{A_{3}}\,\Delta_{k}\overline{u_{3}}(R),$
$A_{1}\mathcal{D}_{k}\Delta_{k}u_{1}(R)+A_{2}\mathcal{D}_{k}\Delta_{k}u_{2}(R)={}$ $A_{3}\mathcal{D}_{k}\Delta_{k}u_{3}(R)+\overline{A_{3}}\,\mathcal{D}_{k}\Delta_{k}\overline{u_{3}}(R),$
where we have defined $u_{1}(r)=J_{k}(\alpha_{+}\,r)$, $u_{2}(r)=J_{k}(\alpha_{-}\,r)$ and $u_{3}(r)=H^{(1)}_{k}(\beta r)$. Since the Helmholtz equation $\Delta_{k}Z_{k}(\mu r)=-\mu^{2}Z_{k}(\mu r)$ is satisfied for $Z_{k}\in\{J_{k},Y_{k},H_{k}^{(1)}\}$, we obtain
$A_{1}u_{1}(R)+A_{2}u_{2}(R)={}$ $A_{3}u_{3}(R)+\overline{A_{3}}\,\overline{u_{3}}(R),$
$A_{1}\mathcal{D}_{k}u_{1}(R)+A_{2}\mathcal{D}_{k}u_{2}(R)={}$ $A_{3}\mathcal{D}_{k}u_{3}(R)+\overline{A_{3}}\,\mathcal{D}_{k}\overline{u_{3}}(R),$
$A_{1}u_{1}(R)-A_{2}u_{2}(R)={}$ $\mathrm{i}\sqrt{\tfrac{1+\tilde{\lambda}}{1-\tilde{\lambda}}}\,\left(A_{3}u_{3}(R)-\overline{A_{3}}\,\overline{u_{3}}(R)\right),$
$A_{1}\mathcal{D}_{k}u_{1}(R)-A_{2}\mathcal{D}_{k}u_{2}(R)={}$ $\mathrm{i}\sqrt{\tfrac{1+\tilde{\lambda}}{1-\tilde{\lambda}}}\,\left(A_{3}\mathcal{D}_{k}u_{3}(R)-\overline{A_{3}}\,\mathcal{D}_{k}\overline{u_{3}}(R)\right),$
which simplifies to
$A_{1}={}$ $\frac{\rho\,\mathrm{e}^{\mathrm{i}\psi}\,A_{3}u_{3}(R)+\rho\,\mathrm{e}^{-\mathrm{i}\psi}\,\overline{A_{3}}\,\overline{u_{3}}(R)}{u_{1}(R)},\hskip 18.49988pt\hskip 18.49988ptA_{2}=\frac{\rho\,\mathrm{e}^{-\mathrm{i}\psi}\,A_{3}u_{3}(R)+\rho\,\mathrm{e}^{\mathrm{i}\psi}\,\overline{A_{3}}\,\overline{u_{3}}(R)}{u_{2}(R)},$
and
$0={}$ $\mathrm{e}^{\mathrm{i}\psi}A_{3}W[u_{1},u_{3}](R)+\mathrm{e}^{-\mathrm{i}\psi}\overline{A_{3}}\,W[u_{1},\overline{u_{3}}](R),$
$0={}$ $\mathrm{e}^{-\mathrm{i}\psi}A_{3}W[u_{2},u_{3}](R)+\mathrm{e}^{\mathrm{i}\psi}\overline{A_{3}}\,W[u_{2},\overline{u_{3}}](R),$
where we recall that
$W[u,v](r)=r\left(u(r)\,v^{\prime}(r)-v(r)\,u^{\prime}(r)\right)=r\left(u(r)\,\mathcal{D}_{k}v(r)-v(r)\,\mathcal{D}_{k}u(r)\right),$
and we have defined
[…]
Since the linear solution is defined up to a constant, we set $A_{3}=\mathrm{e}^{\mathrm{i}\phi}\frac{1}{2\rho\,u_{3}(R)}$ and obtain
[…]
with $\phi$ given by
$\mathrm{e}^{2\mathrm{i}\phi}={}$ $-\mathrm{e}^{-2\mathrm{i}\psi}\frac{W[u_{1},\overline{u_{3}}](R)u_{3}(R)}{W[u_{1},u_{3}](R)\overline{u_{3}}(R)}=-\mathrm{e}^{2\mathrm{i}\psi}\frac{W[u_{2},\overline{u_{3}}](R)u_{3}(R)}{W[u_{2},u_{3}](R)\overline{u_{3}}(R)},$
and where $R,\varepsilon,\tilde{\lambda}$ satisfy
[…]
$={}$ $\mathrm{Im}\left[\mathrm{e}^{\mathrm{i}2\psi}\,W[u_{1},u_{3}](R)\,W[u_{2},\overline{u_{3}}](R)\right]$
$={}$ $0.$
It follows from explicit calculations that
$\begin{split}\mathrm{e}^{\mathrm{i}2\psi}={}&\frac{1}{2}\left(\sqrt{1-\tilde{\lambda}}+\mathrm{i}\,\sqrt{1+\tilde{\lambda}}\right)^{2}=\mathrm{i}\left(\sqrt{1-\tilde{\lambda}^{2}}+\mathrm{i}\,\tilde{\lambda}\right)=\mathrm{i}\,\mathrm{e}^{\mathrm{i}\sin^{-1}(\tilde{\lambda})}\end{split}$
and so
[…]